# Racing Thoughts: Explaining Large Language Model Contextualization Errors

### v1

@@ 5 Related Work
This work takes inspiration from the burgeoning field of mechanistic interpretability, which attempts to understand the circuits and algorithms that a neural network implements in order to accomplish a particular task (Olah, 2022). One fundamental technique in this field is patching, replacing one intermediate representation in an inference pass with another in order to make causal claims about the role of such intermediate representations (Wang et al., 2022; Geiger et al., 2021; Heimersheim & Nanda, 2024). Our work contributes to a trend of using patching (and other techniques in mechanistic interpretability) to study algorithms implemented in production-level LLMs (Ghandeharioun et al., 2024a; Biran et al., 2024; Wu et al., 2024; Lieberum et al., 2023; Feng & Steinhardt, ).
Several recent works have attempted to understand in-context learning (ICL), which might be viewed as an extreme case of context-dependent generation (e.g., Akyürek et al., 2022; Von Oswald et al., 2023; Wies et al., 2024; Falck et al., ; Hahn & Goyal, 2023; Xie et al., 2022). Our work complements this line of research by studying contextualization in more natural settings. Finally, this work contributes to a longstanding body of empirical work identifying distinct processing stages that occur throughout the layers of language models (Tenney et al., 2019). Recent mechanistic work has localized refusal mechanisms to the early layers of LLMs (Ghandeharioun et al., 2024b) and suggested the existence of universal stages of processing (Lad et al., 2024). Earlier work in interpretability has studied the impact of training objective on processing stages (Voita et al., 2019). The present study demonstrates distinct stages of contextualization over layers and suggests interventions that are informed by these stages.
@@ 6 Conclusion
In this work, we introduced a hypothesis to explain contextualization errors in LLMs: the LLM Race Conditions Hypothesis. After establishing a setting for studying contextualization errors systematically, we tested two independent predictions of the LLM Race Conditions Hypothesis: 1) that there exists a critical window when the question tokens are contextualized, and 2) that model failures are caused by improper contextualization of the subject entity during this window. This problem is a natural consequence of purely feedforward architectures, as early layers simply do not have access to later layer representations.
@@ Acknowledgments
The authors would like to thank Sjoerd van Steenkiste, Lucas Dixon, and Apoorv Khandelwal for their helpful discussion and comments on the manuscript.
Appendix A Dataset Partitions
@@
Appendix B Disaggregated Logit Lens Results
@@
Appendix C Autoscoring Details
@@
Consider the following description: GENERATION
Is this description referring to {CORRECT_SENSE, INCORRECT_SENSE}?
Please answer with yes or no:
@@
Appendix D Qualitative Examples of Open-Ended Patchscopes
@@
Appendix E Pairwise Patching Results
@@
Appendix F Replicating results on Llama-2
In this section, we replicate the behavioral analysis found in § 3 and causal interventions found in § 4.1 and 4.2 on Llama-2-13b-chat-hf (Touvron et al., 2023). In Fig. 10(a) and Fig. 10(b), we find the same trends as observed in § 3: performance degrades as the number of distractors increase, and performance depends on cue position. However, we observe an even more reliable U-shaped curve with cue position using Llama-2.
@@ Appendix F Replicating results on Llama-2
(a) Accuracy vs. number of distractors in Llama-2. We find that injecting distractor text into the prompt causes sharp performance degredations across all datasets.
(b) Performance on the 5-distractors dataset, disagreggated by the position of the cue amongst distractor sentences. We reliably find evidence of a U-shaped performance curve.
Figure 10: Behavioral results across all three datasets for Llama-2. We find that injecting distractors routinely engenders contextualization errors.

### v2

@@ 5 Related Work
This work takes inspiration from the burgeoning field of mechanistic interpretability, which attempts to understand the circuits and algorithms that a neural network implements in order to accomplish a particular task (Olah, 2022). Our work contributes to a trend of using patching to study algorithms implemented in production-level LLMs (Ghandeharioun et al., 2024a; Biran et al., 2024; Wu et al., 2024; Lieberum et al., 2023; Wang et al., 2022; Geiger et al., 2021; Heimersheim and Nanda, 2024).
Several recent works have attempted to understand in-context learning (ICL), which might be viewed as an extreme case of context-dependent generation (e.g., Akyürek et al., 2022; Von Oswald et al., 2023; Wies et al., 2024; Falck et al., ; Hahn and Goyal, 2023; Xie et al., 2022). Our work complements this line of research by studying contextualization in more natural settings. Additionally, recent work has analyzed knowledge conflicts — contexts in which information encoded within the weights of an LLM conflict with information provided in context (Xu et al., 2024). The Factual dataset is comprised of knowledge conflicts regarding countries and their capitals. Thus, we might hypothesize that race conditions may be present in other contexts that contain knowledge conflicts.
Finally, this work contributes to a body of work that identifies distinct processing stages throughout the layers of language models (Tenney et al., 2019). Recent mechanistic work has localized refusal mechanisms to the early layers of LLMs (Ghandeharioun et al., 2024b) and suggested the existence of universal stages of processing (Lad et al., 2024). Earlier work in interpretability has studied the impact of training objectives on processing stages (Voita et al., 2019). The present study demonstrates distinct stages of contextualization over layers and suggests interventions that are informed by these stages.
@@ 6 Conclusion
In this work, we introduced a hypothesis to explain contextualization errors in LLMs: the LLM Race Conditions Hypothesis. After establishing a setting for studying contextualization errors systematically, we tested two predictions of the LLM Race Conditions Hypothesis: 1) that there exists a critical window when the question tokens are contextualized, and 2) that model failures are caused by improper contextualization of the subject entity during this window. This is a natural consequence of purely feedforward architectures, as early layers simply do not have access to later layer representations.
@@
7 Limitations
This work conducts a mechanistic analysis of one form of contextualization error that occurs in LLMs. Though the LLM Race Conditions Hypothesis appears to apply across a range of settings (as indicated by the diverse datasets employed in this work), it does not comprehensively explain all LLM contextualization errors. In particular, errors due to missing factual knowledge (Hase et al., 2023) lie outside the scope of this work. Additionally, this work suggests several remediations, including those based on the inference-time interventions presented in § 4.2. However, we note that these interventions are not practical in their current state, as they require one to know which entities to patch and which layers to use as source and target a priori. Future work must develop these ideas further before employing them in a production system to address LLM Race Conditions.
8 Ethical Considerations
This work demonstrates that contextualization errors occur in settings where a model is likely to encode a gender bias associated with a particular profession. This adds a small contribution to a longstanding literature describing gender biases in language models (May et al., 2019; Rudinger et al., 2018; Weidinger et al., 2021). As noted in § 7, none of the interventions presented here should be used to address issues of gender bias in production language models.
@@ Acknowledgments
The authors would like to thank Sjoerd van Steenkiste, Lucas Dixon, and Apoorv Khandelwal for their helpful discussion and comments on the manuscript and Rachel Goepner for proofreading the manuscript.
Appendix A Serial Position Effects
Figure 8: Performance on the 5-distractors dataset, disagreggated by the position of the cue amongst distractor sentences.
Liu et al. (2024) show serial position effects—multi-document question-answering performance is highest when the relevant document is presented at the beginning or at the end of a set of documents, and lowest when it occurs in the middle. However, their work studies contexts that contain thousands of tokens, whereas our longest questions are on the order of 100-200 tokens. Interestingly, we also observe sensitivity to the position of relevant information: Fig. 8 presents results from our five-distractor datasets, indexed by the position of the cue amongst the distractors. The model often performs better when the cue occurs before the distractors, and sometimes (i.e., in the Polysemous dataset) demonstrates the U-shaped curve found in Liu et al. (2024).
Appendix B Dataset Partitions
@@
We provide three examples from each of the datasets used to analyze gemma-2-9b-it. Subject entities are in yellow, cues are in blue, questions are in green, and answers are provided in bold for the reader (though these are not included in the prompt provided to the model).
Polysemous Word Dataset
  1. 1.
    Please answer succinctly. As governor Beckham sought to unite his party and the state . Past this interchange , Route 50 widens to a four @-@ lane divided highway and reaches a full interchange with the Atlantic City Expressway . I am in an orchard. Mr. I see an apple. Is it a fruit? Answer: Yes
  2. 2.
    Please answer succinctly. On 21 March 1918 the German Army launched its Spring Offensive in a last @-@ ditched effort to win the war , unleashing sixty @-@ three divisions over a 70 miles ( 110 km ) front . Altar 10 is carved with a captive tied to a scaffold . I am in a living room. The era of the wooden steam ship @-@ of @-@ the @-@ line was brief , because of new , more powerful naval guns . I look at the chair. Is it a piece of furniture? Answer: Yes
  3. 3.
    "Please answer succinctly. Family Edopidae . Churchill , Ward ( Winter – Spring 2003 ) . I am at the zoo. On April 3 , Brad Stevens and the Butler Bulldogs faced off with Michigan State in the national semi @-@ finals . I see a seal. Is it an emblem? Answer: No
Factual Dataset
  1. 1.
    Answer based on the information provided here. Sylvester detects the influence of late works by Titian in other aspects , especially in the deep and rich colouring , Velázquez ’s portrayals of Philip IV , and agrees with identification of pastels of Edgar Degas as a source . Forget everything you know about geography. The capital city of Austria was just renamed from Vienna to Grytviken. Is the capital city of Austria named Vienna? Answer: No
  2. 2.
    Answer based on the information provided here. Waldman felt that " TNA put on a strong PPV " that was in " large part to little gimmickery . Forget everything you know about geography. The capital city of Bolivia was just renamed from Sucre to Willemstad. Is the capital city of Bolivia named Sucre? Answer: No
  3. 3.
    Answer based on the information provided here. George Vincent 1920 @-@ 24 . Forget everything you know about geography. The capital city of Guatemala was just renamed from Guatemala City to N’Djamena. Is the capital city of Guatemala named N’Djamena? Answer: Yes
Gender Dataset
  1. 1.
    Please answer succinctly. The CEO is somebody’s grandmother. As the nominations for the 72nd Academy Awards approached , a frontrunner had not emerged . Is the CEO a woman? Answer: Yes
  2. 2.
    Please answer succinctly. The baker is somebody’s grandmother. Ma Anand Sheela served 29 months in a minimum security federal prison for charges related to assault , attempted murder , arson , wiretapping and the 1984 bioterror attack in The Dalles , and moved to Switzerland after her release from prison in 1988 . Is the baker a man? Answer: No
  3. 3.
    Please answer succinctly. The mechanic is somebody’s grandfather. US 40 and NJ 50 follow the Harding Highway , a two @-@ lane undivided road , turning to the northeast and crossing the Great Egg Harbor River . Is the mechanic a man? Answer: Yes
Appendix C Disaggregated Logit Lens Results
@@
Appendix D Autoscoring Details
@@
Consider the following description: {GENERATION} Is this description referring to {CORRECT_SENSE, INCORRECT_SENSE}? Please answer with yes or no:
@@
Appendix E Qualitative Examples of Open-Ended Patchscopes
@@
Appendix F Pairwise Patching Results
@@
Appendix G Replicating results on Llama-2
In this section, we replicate the behavioral analysis found in § 3 and causal interventions found in § 4.1 and 4.2 on Llama-2-13b-chat-hf (Touvron et al., 2023). In Fig. 11 and Fig. 12, we find the same trends as observed in § 3: performance degrades as the number of distractors increase, and performance depends on cue position. However, we observe an even more reliable U-shaped curve with cue position using Llama-2.
@@ Appendix G Replicating results on Llama-2
Figure 11: Accuracy vs. number of distractors in Llama-2. We find that injecting distractor text into the prompt causes sharp performance degradations across all datasets.
Figure 12: Performance on the 5-distractors dataset in Llama-2, disagreggated by the position of the cue amongst distractor sentences. We reliably find evidence of a U-shaped performance curve.
@@
Appendix H Replicating results on Gemma-2-2b
In this section, we replicate all of our main results using Gemma-2-2b-it, a smaller (i.e., 26 layer) version of Gemma-2-9b, on the Polysemous word dataset. Overall, we find the same pattern of results that was identified in the larger model. Behavioral results are displayed in Figure 15. We identify the data partition with 3 distractors and cue index 2 to run our intervention analyses on, as this partition results in 49.2% accuracy.
Attention mass analyses are displayed in Figure 16. Here, we display the summed mass allocated to the subject entity across attention heads in Subfigure 16(a) and the max attention mass allocated to the subject entity from any individual head across layers in Subfigure 16(b). The patterns are largely the same across both graphs, both reflecting an increase in attention mass in the middle (approximately 7-18) layers. This indicates that the attention mass patterns described in the main text also occur at the level of individual attention heads.
Logit lens results are displayed in Figure 17, and attention interventions are displayed in Figure 18. Interestingly, both of these analyses indicate that the critical window ends around layer 10, near the midpoint of the model layers. Notably, the critical window of Gemma-2-9b also ends near the midpoint of that model’s layers (around layer 20). This suggests that the critical window may shrink as the number of model layers shrink, rather than always ending after a specific amount of processing. If this hypothesis was verified, it could provide a partial algorithmic account for the emergence of impressive contextualization capabilities in deeper LLMs.
Finally, we provide the results of patching interventions in Figure 19.
Figure 15: Accuracy vs. number of distractors in Gemma-2-2b. We find that injecting distractor text into the prompt causes performance degradations.
Figure 16: (a) Attention mass analysis on Gemma-2-2b generated by summing attention mass across all heads. (b) Attention mass analysis on Gemma-2-2b generated by taking the max attention mass from a single head at each layer.
Figure 17: Replicating the logit lens analysis in Gemma-2-2b. We find evidence that the critical window ends around layer 10.
Figure 18: Replicating the attention ablation intervention in Gemma-2-2b. We find evidence that the critical window ends around layer 10.
Figure 19: Replicating the patching interventions in Gemma-2-2b.